Find the extension produced in a copper wire of length 2 m and diameter 3 mm when a load of 30 N is applied. (Young’s modulus for copper = 11 x 10^11 Nm-2)
1
Expert's answer
2020-10-29T07:03:12-0400

Mechanical stress:

"\sigma=\frac{F}{S}=\epsilon E"

where "F" - is a force of the load.

"S=\frac{\pi d^2}{4}"

"\epsilon=\frac{\Delta l}{L_0}"

"E" - Young's module.


We can rewrite this formula as

"\frac{4F}{\pi d^2}=\frac{\Delta l}{L_0}E"

hence

"\Delta l=\frac{4FE}{\pi d^2}l_0"